Dundee: Dom, Simon, Petr, Jason, Riad, June,
Remote: Mark, Josh, Will, Frances, J-m, Melissa, Liza, Chris, Emil, David
Start: 2:00 pm
(2-3 minutes each)
Prod63: HPA run 03 images still importing.
idr0050 (Springer), reorganising data to improve visualisation
modified this week for release
search (tests failing), other tests passing.
More microservice work
Planning some Bio-Formats work
Call this morning regarding ongoing RDF work (see gitlab)
need to link RDF terms with OME-XML terms on a ome/design issue
zarr-specs repository with language agnostic design
upcoming with HDF (28th) and HCA (~5th) engineers
Several companies taking part in discussions
French meeting in progress
User/training meetings (Jason):
(5 min. max; tech. Discussion should be highlighted to relevant people and rescheduled)
old demo server users before deleting their data? Could delete just underlying image files but inconsistently leave objects in database intact.
JRS: number of logins
improve after we start adding public data)
Get cleaned up
Identify ongoing (anonymized) queries
(20-25 minutes plus 15 minutes questions max)